# The Riemann Hypothesis is false

### v36

@@
Suppose that $\Theta<1$. Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
uniformly for $-1\leq\sigma\leq 2$. Let $n\geq\frac{T}{\log^{2}T}$. By the residue theorem, we have
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $T\leq t\leq H$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
$\zeta(\sigma+it)=\sum_{n\leq x}n^{-\sigma-it}+\chi(\sigma+it)\sum_{v\leq y}v^{\sigma-1+it}+O(x^{-\sigma}\log t+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h$ is a constant and $|\chi(\sigma+it)|=\Big{|}\frac{\zeta(\sigma+it)}{\zeta(1-\sigma-it)}\Big{|}=\Big{|}\Big{(}\frac{2\pi}{t}\Big{)}^{\sigma-\frac{1}{2}+it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{1/2-\sigma}$ [6, p.78]. Thus since $|\zeta(it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows from (4) that
$S(x):=\sum_{n\leq x}n^{it}\ll T^{1/2}\log^{3}T$ (5)
uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}$. Then by partial summation, we obtain from (5) that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
for $\sigma\in\mathbb{R},|\sigma|\leq 2$. Combining (1) and (6) yields $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$ (see e.g [3, p.403]), note that
@@
for $\sigma\geq\sigma_{0}>0.$ Combining (6) and (7) gives $\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{3}\ll T^{1/2}\log^{8}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{8}T.$ (8)
For $\Re(s)>1$, note that $-\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$. Putting this into $I_{1}$ gives
$I_{1}=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{e\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{e\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}<n<T$ (unless specified otherwise). For $2\leq m\leq T(1-\varepsilon)$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-\frac{1}{\log n}-1}(\log(m)\log(n)-\log^{2}(m)-\log^{2}n)}{n\log^{2}(n)\log^{2}(n/m)}$. Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) by routine partial summation that
$\sum_{2\leq n\leq m(1-\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
@@
uniformly for $m\geq T(1+\varepsilon)$ and $2\leq m\leq T(1-\varepsilon)$. For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll T^{1/4}$ so it similarly follows from (5) by partial summation that
$\sum_{\frac{T}{\log^{2}T}<n\leq\frac{m}{1+\varepsilon}<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
@@
$\sum_{\frac{T}{\log^{2}T}<m(1+\varepsilon)\leq n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
We now treat the ranges $\frac{m}{1+\varepsilon}<n\leq m-1$ and $m+1\leq n<m(1+\varepsilon)$. For $m\ll T$, note that
$\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\log T$ (15)
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\log T$ (16)
Notice that (16) is obtained by writing $m=n+r$ for $1\leq r\leq n\varepsilon$ and applying the fact that $\log(1+\delta)\sim\delta$ uniformly for $1/m\leq\delta\leq\varepsilon$, hence
$\sum_{\frac{n}{1+\varepsilon}\leq m<n-1}\frac{1}{m\log(n/m)}\ll\sum_{1\leq r\leq\varepsilon m-1}\frac{1}{r}\ll\log T.$ (17)
A similar argument yields (15). Define $\psi(x):=\sum_{m\leq x}\Lambda(m)$. Note that $\psi(x)=x+O(x^{\Theta}\log^{2}x)$ [5, p.430]. Thus from (16) it follows that
$\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $\ll\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (18)
$\ll\log T\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\Lambda(m)$ (19)
$=(\log T)(\psi(T+T\varepsilon)-\psi(T-T\varepsilon)+O(\log T))$ (20)
$=(\log T)(2\varepsilon T+O(T^{\Theta}\log^{2}T))$ (21)
$=o(T).$ (22)
By arguing similarly, one obtains from (15) that
$\sum_{T(1-\varepsilon)<m<T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (23)
Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation and invoking (23), (22), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (24)
@@
since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Theorem 1 of [3], we have
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (29)
@@
By Theorem 4.11 of [6], we have
$\sum_{1\leq n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (31)
uniformly for $T\leq\gamma_{T}\leq H.$ Thus
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (32)
$\ll T^{\Theta}\log T.$ (33)
@@
(i) From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (26) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$. Note that the analogue of (26) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
(ii) The key parts of our argument, are the estimates in (12), (13), (14), (22) and (23). The first three are are routine applications of partial summation from (5), whilst (22) and (23) follow easily from the well-known estimates for $\psi(x)$.

### v37

@@
Lemma 1. (Theorem 4.11 of [6]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0,|\Im(s)|<2\pi x/C$, when $C$ is a given constant $>1$.
Lemma 2, (Theorem 1 of [3]). One has $\sum_{T\leq\gamma_{T}\leq 2T+1}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ uniformly for $2\leq n\ll T.$
@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+\varepsilon+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+\varepsilon+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{\Theta+\varepsilon+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
say. From now on, let $T\leq t\leq H$ and $\sigma=\Re(s)$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [6]), note that
$\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [6, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (5) that
$S(x):=\sum_{n\leq x}n^{it}\ll T^{\frac{1}{2}\log^{3}T}$ (5)
uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}.$ Then by Abel summation, it follows that
$\sum_{n\leq x}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$, we have
@@
for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{7}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)=\sigma>1$. Inserting this into $I_{1}$ gives
$I_{1}=\frac{(T-H))(\Lambda(n))}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$, $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
@@
uniformly for $1\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (13)
@@
$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=m/r(1+O(r/m))$ hence
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{it}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(r/m^{2})\Big{)}$ (16)
$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
If $f(x)$ is a real function that is differentiable thrice and $f^{\prime}(x)$ is steadily decreasing on $[a,b]$, note that
$\sum_{a<n\leq b}g(n)e^{2\pi if(n)}\sim e^{-\frac{1}{4}\pi i}\sum_{\alpha<v\leq\beta}\frac{e^{2\pi i(f(x_{v})-vx_{v})}}{|f^{\prime\prime}(x_{v})|^{\frac{1}{2}}}g(x_{v})\ll\sum_{\alpha<v\leq\beta}\Big{|}\frac{g(x_{v})}{f^{\prime\prime}(x_{v})^{\frac{1}{2}}}\Big{|}$ (18)
where $f^{\prime}(x_{v})=v,\alpha=f^{\prime}(a)$ and $\beta=f^{\prime}(b)$ [6, p.78]. Taking $g(u)=1/u,f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},f^{\prime}(u)=\frac{t}{2\pi(m+u)},f^{\prime\prime}(u)=-\frac{t}{2\pi(m+u)^{2}},x_{v}=\frac{t}{2\pi v}-m,f^{\prime\prime}(x_{v})=\frac{v^{2}}{t}$ gives, for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$,
$\sum_{0<r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}=\sum_{0<r\leq\varepsilon m-1}\frac{e^{it\log(1+\frac{r}{m})}}{r}\ll\sum_{\alpha<v\leq\frac{t}{2\pi m(1+\varepsilon)}}\frac{\sqrt{t}}{\frac{t}{2\pi}-mv}\ll\sum_{\alpha<v\leq\log^{2}t}\frac{1}{\varepsilon\sqrt{t}}\ll T^{-1/4}\log^{2}T.$ (19)
Inserting (19) into the right-hand side of (17), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (20)
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-m/r(1+O(r/m))$. Thus by arguing as above, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll T^{3/4}\log^{2}T.$ (21)
Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (21), (20), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi n}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (22)
@@
since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Lemma 2, we have
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n<T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (27)
@@
Suppose that $\Theta<1$. Then by Lemma 1 we have
$\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (29)
uniformly for $T\leq\gamma_{T}\leq H.$ Hence
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}\ll T^{\Theta}\log T.$ (30)
@@
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (24) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$. Note that the analogue of (21) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.